1) Helium is commonly used to fill balloons for parties. How many 2.50 L balloons can be filled at SATP with a container that contains 101g of helium? *


2) You are learning about Charle's Law and Boyle's Law. Describe BOTH laws using the kinetic molecular theory (KMT) explain their relationships, and make sure to provide an equation for each law


1
Expert's answer
2020-06-29T07:52:10-0400

1) pV=nRT

n= m/Mr = 101/4 = 25.25

V= nRT/p = 25.25 x 8.314 x 273 / 101.325 = 565,610466 L.

Ballons = 565,610466/2.5 = 226.

2) Kinetic Molecular Theory states that gas particles are in constant motion and exhibit perfectly elastic collisions. Kinetic Molecular Theory can be used to explain both Charles' and Boyle's Laws. The average kinetic energy of a collection of gas particles is directly proportional to absolute temperature only.